# Artificial Neural Networks for Magnetoencephalography: A review of an emerging field

### v1

@@ 2.3 Notes on terminology
Some of the ML nomenclature overlaps with terminology used in M/EEG research, which can sometimes lead to confusion. In particular, the term ’epoch’ in deep learning generally describes a complete pass of the training data. Meanwhile, in the field of M/EEG, it usually describes a segment of data (or a single trial). In addition, the term ’samples’ in ML usually refers to individual examples from the dataset, which should not be confused with the use of the same term to refer to a data point in a M/EEG time series, where the parameter known as ’sampling rate’ (Hz) refers to the rate at which the data is assessed, i.e., data points per second.
@@ 3 Results
The methodology explained in the previous section (2) led to the identification of 119 relevant studies. Among the papers considered, 102 were peer-reviewed, while 17 had not yet undergone peer review at the time of writing. Our categorization method resulted in the distribution of 69 papers under ’Classification,’ accounting for 58% of the total; 16 papers were categorized as ’Modeling,’ making up 13.4%; and 34 papers fell into the ’Other’ category, representing 28.6% of the study corpus included in this review (See Fig 1). The category containing most of the papers is the ’Classification’ category, followed by ’Other’; the category regrouping the least amount of studies is ’Modeling’. Generally, the trend of using ANN on MEG data has increased over the years. A similar growth trajectory is anticipated for studies on this topic, analogous to the rapid expansion of research utilizing Artificial Neural Networks (ANNs) for Electroencephalography (EEG) data analysis ([17]).
3.1 General Overview
The in-depth survey revealed several key observations about this emerging field (Fig 1). Interestingly, out of the 119 selected studies for this review, 64 use less than 20 subjects, 34 use between 20 and 100 subjects, and 15 use more than 100 subjects. The remaining studies use synthetic MEG data or do not mention the number of subjects, or patients included in the study. The number of samples (number of trials or epochs) in datasets varies across a wide range, from 17 to 150000, with an average of 76261, a standard deviation of 289279, and a median of 3540. However, a few outliers significantly impact the average and standard deviation of the number of samples used in the included studies (Correcting for outliers gives an average sample number of 13249 and a standard deviation of 21247). Furthermore, the length of the data segments fed to the neural network varies between 300ms to 10s. The most commonly used sampling frequencies are 1000Hz (26 papers) and 250Hz (16 papers), and the median across all studies was 600Hz.
@@ 3.1 General Overview
Category Subcategory Study
Classification Decoding [18], [19], [20], [21], [22], [23], [24], [25], [26], [27],
[28], [29], [30], [31], [32], [33], [34], [35], [36],[37],
[38], [39], [40], [41], [42], [43], [44], [45]
BCI [46], [47], [48], [49], [50], [51], [52], [53],
[54], [55], [56], [57], [58], [59]
Clinical [60], [61], [62], [63], [64], [65], [66], [67], [68], [69],
[70], [71], [72], [73], [74], [75], [76]
Event detection [77], [78], [79], [80], [81], [82], [83], [84], [85][86]
Modeling Visual Cortex [87], [88], [89], [90], [91], [92], [93], [94], [95] [96], [97]
Auditory Cortex [98], [99], [100], [101], [102], [103]
Other Preprocessing [104], [105], [106], [107], [108], [109], [110], [111],
Methods [112], [113], [114], [115], [116], [117], [118], [119],
[120], [121], [122], [123], [124], [125], [126], [127]
Source Localization [128], [129], [130], [131], [132], [133], [134],
[135], [125]
2016 Yu et al Decoding 15 2000 CNN (3) train/test split
2017 Wang et al Decoding 2 410 MLP (1)
4Fold on train +
holdout test set
2018 Hramov et al Decoding 5 150000 MLP (1) train/valid split
2018
Frolov
and Pisarchik
Decoding 5 10000 MLP (1) train/test split
2019 Dash et al Decoding 4 N/M
MLP (3)
LSTM
train/valid/test split
2019 Garry et al Decoding 605 75396 CNN (3) train/valid/test split
2019 Kim et al Decoding 27 5400 CNN + Attention (6) 5Fold
2019 Kostas et al Decoding 101 3900 CNN and RCNN (N/A) 5Fold
2019 Dash et al Decoding 8 2400 MLP (1) train/test split
2019 Huang and Yu Decoding 16 588 CNN (4) train/test split
2020 Abdellaoui et al Decoding 18 N/M CNN + Attention + LSTM (10) train/test split
2020 Dash et al Decoding 7 300 MLP (1) 5Fold
2020 Dash et al Decoding 10 180 MLP (1) 5Fold
2021 Li et al Decoding 16 9414 2 x GRU + FC LOO
2021 Feng et al Decoding 200 397 GoogLeNet - CNN (144) train/valid split
2021 Chang et al Decoding 8 108 GAN-like (5) Holdout set
2021 Pilyugina et al Decoding 17 17 CNN (1) train/test split
2022 Engemann et al Decoding 646 N/M ShallowFBCSPNet + Deep4Net 5Fold
2022 Csaky et al Decoding 15 53100 WaveNet* train/valid split
2022 Shi et al Decoding 17 10880 EEGNet (4) nested LOSO
2023 Zhang et al Decoding 20 N/M N/M Leave One Out
2023 Csaky et al Decoding 15 53100 Wavenet Classifier (9) LOSO
2023 Özer et al Decoding 18 9414 LSTM, GRU and CNN 10Fold
2023 Bu et al Decoding 12 8640 CNN (4) 5Fold + train/test split
2024 Boyko et al Decoding 96 2304000 N/M train/valid/test split
2024 Yang et al Decoding 27 179977 N/A train/valid/test split
2024 Zubarev et al Decoding 12 3600 LF-CNN (3) 9Fold
2024 Yang et al. Decoding 27 179977 NeuGPT train/valid/test
2018 Meng et al Clinical 45 N/M CNN (4) 4Fold
2019 aoe et al Clinical 233 70000 CNN (12) 10Fold
2020 Gu et al Clinical 32 N/M MLP (4) + Self-attention head cross-validation
2020 Zhang et al Clinical 32 N/M LSTM (2) 10Fold + holdout test set
2021 Xu et al Clinical 129 N/M G2G network (1) train/valid split
2021 Giovannetti et al Clinical 87 6525 AlexNet - CNN (5) LOSO
2021 Wu et al Clinical 38 2280 CNN (4) 2Fold
2021 Huang et al Clinical 95 95 EEGNet (4) train/valid/test split
2022 Huang et al Clinical 190 33500 ResNet (5) 5Fold + train/val/test split
2022 Fujita et al Clinical 180 21000
MLP (1)
MNet (12)
10Fold + holdout test set
2023 Barik et al Clinical 60 N/M MLP (1) 5Fold nested
2024 Anand et al Clinical N/M 1080000 N/M train/test split
2024 Achterberg et al Clinical 8 2400 DNN (4) train/valid/test split
2018 Dash et al BCI 4 1635 MLP (1) train/valid/test split
2018 Dash et al BCI 3 1225 CNN (8) 5Fold (trials)
2019 Hramov et al BCI 10 N/M MLP N/M
2019 Hramov et al BCI 7 N/M MLP (2) N/M
2019 Zubarev et al BCI 7 11354 LF-CNN (3) LOO
2019 Dash et al BCI 4 1635 CNN (5) train/valid/test split
2020 Dash et al BCI 8 3046
AlexNet (7)
ResNet101 (101)
Inception-ResNet-v2 (164)
train/valid/test split
2020 Dash et al BCI 8 3046 LSTM (3) N/M
2020 Yeom et al BCI 9 2160 LSTM 5Fold
2020 Dash et al BCI 4 1500 LSTM* train/valid/test split
2020 Lopopolo et van den Bosch BCI 15 1377 CNN (4) houldout test set
2021 Ovchinnikova et al BCI 32 422 LF-CNN (3) nested 4Fold in a 5Fold
2022 Fan et al BCI 61 39040 CNN (3) train/test split
2018 Guo et al Event 10 102 SSAE (3) 5Fold
2019 Zheng et al Event 20 4000 CNN (8) multiple KFold
2020 Liu et al Event 20 150 CNN* KFold
2021 Zhang Chuncheng et al Event 10 N/M EEGNet (4) cross-validation
2022 Hira et al Event 348 23177 ResNet (26) and AE 5Fold and 10Fold
2022 Bhat et al Event 15 11000 N/M 5Fold
2022 Zhao et al Event 20 1320 DANN Leave One Out
2022 Guo et al Event 20 202 Transformer KFold
2022 Zhang et al Event 20 150
CADNet
DendriteNet
N/M
2023 Zheng et al Event 48 N/M CNN (5) 7Fold
2023 Mouches et al Event 95 40662 N/M 10Fold
2024 Wei et al Event 277 230325 N/M train/valid split
2024 He et al Event 11 35013 CNN + Attention (5) LOSO
2024 Dev et al Event 7 7060 N/M train/valid/test split
2024 Hira et al. Event 1782 60139 26 Layer scSE-ResNet 5Fold subject-wise
2016 Cichy et al Visual 15 3540 SuperVision - CNN () N/M
2017 Cichy et al Visual 15 N/M CNN (8)
Bootsrapping
resampling
2017 Seeliger et al Visual 15 1000 VGG-S - CNN (10) 10Fold
2018 Dima et al Visual 19 720
CNN (7)
AlexNet - CNN (8)
5Fold
2018 Bankson et al Visual 32 32 VGG-F - CNN (7) LOSO
2019 Rajaei et al Visual 15 1536
AlexNet - CNN (7)
HRRN (152)
LOO
2019 Kietzmann et al Visual 15 380
CNN
RCNN (6)
2Fold
bootstrapping
2020 Giari et al Visual 25 3525 AlexNet - CNN (5) N/A
2022 van Vliet et al Words 15 560 VGG (11) N/A
2023 von Seth et al Visual 36 N/M N/M N/M
2019 Donhauser Speech 11 77 LSTM* 7Fold
2022 Caucheteux and King Auditory 92 N/M
Transformer
CNN (4, 8 or 12)
N/A
2022 Wingfield et al Speech 16 66300 MLP (6) N/A
2023 Desbordes et al Speech 11 3630
LSTM (2)
Transformer (11)
10Fold
2024 Brodbeck et al
Speech
Recognition
18 N/M N/M N/M
2024 Lyu et al
Speech
Recognition
16 5760 N/M N/M
2019 Dinh et al SL 1 N/M LSTM train/valid split
2021
Pantazis
and Adler
SL N/A N/M CNN (5) and MLP (4) N/A
2021 Dinh et al SL 1 1653 LSTM train/valid split
2022 Sun et al SL 29 620256 ResNet (4) LSTM (3) train/test split
2023 Sun et al SL 29 620256 N/M N/M
2023 O’Reilly et al SL 1 446 SimpleRNN (4) Bootstrapping
2024
Sanchez-Bort
et al
SL 1 N/M MPSS (4) N/M
2024 Jiao et al SL 1 N/M CNN + Attention N/M
2024 Yokoyama et al. SL 3 N/M 4LCNN (4) train/valid sets
2016
Hyvärinen
and Morioka
Preprocessing 9 N/M FC (4) M/M
2017 Garg et al Preprocessing 49 980 CNN (6) train/test split
2017 Garg et al Preprocessing 44 880 CNN (9) LOO + train/test split
2018 Croce et al Preprocessing 67 4038
CNN (6); FC (6)
CNN+FC
10Fold (trials)
2018 Hasasneh Preprocessing 48 1112
temporal CNN (3)
spatial CNN (2)
sCNN + tCNN
50Fold
2021 Feng et al Preprocessing 4 66780 GoogLeNet - CNN (144) train/valid/test split
2021 Treacher et al Preprocessing 217 294 CNN (11) 10Fold
2023 Hamdan et al Preprocessing N/M N/M AE (9) N/M
2017 Guo et al Method 3 N/M CNN (5) 10Fold
2019 Harper et al Method 72 N/M LRCN train/valid/test split
2020 Shu and Fyshe Method 9 60 SAE LOO
2021 Abdellaoui et al Method 18 N/M CNN + Attention train/test split
2022 Priya and Jayalakshmy Method 23 13472 GoogLeNet
KFold with
holdout test set
2023 Gosti et al Method 10 N/M RHoMM - RNN (N/M) N/M
2023 Elshafei et al Method 1 166800
DNN (4)
CNN (1)
train/test split
2023 Fan et al Method 669 70000
LFCNN
VARCNN
HGRN
train/valid split
2023 Csaky et al Method 37 53100 WaveNet-based train/valid split
2023 Zhu et al Method 676 100000
MLP (1)
NICA(TCL) (3)
NICA(IIA) (3)
N/A
2024 Solana et al Method 16 N/M
ROCKET-based
models (3)
train/valid split
2024 Fan et al Method 250 29238
EEGNet
ResNet
ShuffleNet
N/M
2024 Gallard et al Method 44 N/M CycleGAN N/A
2024 Chou et al Method 17 N/M CNN (3) train/valid/test split
2024 Csaky et al Method 15 3540 GPT train/valid/test split
2024 Gideoni et al. Method 6 N/M MLP and CNN train/valid/test split
2024 Ferrante et al. Method 4 N/M CNN and CLIP N/A

### v3

@@ 2.1 Litterature research
2016 Yu et al. [49] Decoding 15 2000 CNN (3) train/test split
2017 Wang et al. [50] Decoding 2 410 MLP (1)
4Fold on train +
holdout test set
2018 Hramov et al. [51] Decoding 5 150000 MLP (1) train/valid split
2018
Frolov
and Pisarchik [52]
Decoding 5 10000 MLP (1) train/test split
2019 Dash et al[53] Decoding 4 N/M
MLP (3)
LSTM
train/valid/test split
2019 Garry et al. [54] Decoding 605 75396 CNN (3) train/valid/test split
2019 Kim et al. [55] Decoding 27 5400 CNN + Attention (6) 5Fold
2019 Kostas et al. [56] Decoding 101 3900 CNN and RCNN (N/A) 5Fold
2019 Dash et al. [57] Decoding 8 2400 MLP (1) train/test split
2019 Huang and Yu [58] Decoding 16 588 CNN (4) train/test split
2020 Abdellaoui et al. [59] Decoding 18 N/M CNN + Attention + LSTM (10) train/test split
2021 Dash et al. [60] Decoding 7 300 MLP (1) 5Fold
2020 Dash et al. [61] Decoding 10 180 MLP (1) 5Fold
2021 Li et al. [62] Decoding 16 9414 2 x GRU + FC LOO
2021 Feng et al. [63] Decoding 200 397 GoogLeNet - CNN (144) train/valid split
2021 Chang et al. [64] Decoding 8 108 GAN-like (5) Holdout set
2021 Pilyugina et al. [65] Decoding 17 17 CNN (1) train/test split
2022 Engemann et al. [66] Decoding 646 N/M ShallowFBCSPNet + Deep4Net 5Fold
2021 Shi et al. [67] Decoding 17 10880 EEGNet (4) nested LOSO
2023 Zhang et al. [68] Decoding 20 N/M N/M Leave One Out
2023 Csaky et al. [69] Decoding 15 53100 Wavenet Classifier (9) LOSO
2023 Özer et al. [70] Decoding 18 9414 LSTM, GRU and CNN 10Fold
2023 Bu et al. [71] Decoding 12 8640 CNN (4) 5Fold + train/test split
2024 Boyko et al. [72] Decoding 96 2304000 N/M train/valid/test split
2024 Yang et al. [73] Decoding 27 179977 N/A train/valid/test split
2024 Zubarev et al. [74] Decoding 12 3600 LF-CNN (3) 9Fold
2024 Yang et al. [75] Decoding 27 179977 NeuGPT train/valid/test
2024 Jayalath et al. [76] Decoding 900 2296800 CNN + LSTM train/valid/test
2020 Shu and Fyshe [77] Decoding 9 60 SAE LOO
2018 Meng et al. [78] Clinical 45 N/M CNN (4) 4Fold
2019 aoe et al. [79] Clinical 233 70000 CNN (12) 10Fold
2020 Gu et al. [80] Clinical 32 N/M MLP (4) + Self-attention head cross-validation
2020 Zhang et al. [81] Clinical 32 N/M LSTM (2) 10Fold + holdout test set
2021 Xu et al. [82] Clinical 129 N/M G2G network (1) train/valid split
2021 Giovannetti et al. [83] Clinical 87 6525 AlexNet - CNN (5) LOSO
2021 Wu et al. [84] Clinical 38 2280 CNN (4) 2Fold
2021 Huang et al. [85] Clinical 95 95 EEGNet (4) train/valid/test split
2022 Huang et al. [86] Clinical 190 33500 ResNet (5) 5Fold + train/val/test split
2022 Fujita et al. [87] Clinical 180 21000
MLP (1)
MNet (12)
10Fold + holdout test set
2023 Barik et al. [88] Clinical 60 N/M MLP (1) 5Fold nested
2024 Anand et al. [89] Clinical N/M 1080000 N/M train/test split
2024 Achterberg et al. [90] Clinical 8 2400 DNN (4) train/valid/test split
2018 Dash et al. [91] BCI 4 1635 MLP (1) train/valid/test split
2018 Dash et al. [92] BCI 3 1225 CNN (8) 5Fold (trials)
2019 Hramov et al. [93] BCI 10 N/M MLP N/M
2019 Hramov et al. [94] BCI 7 N/M MLP (2) N/M
2019 Zubarev et al. [95] BCI 7 11354 LF-CNN (3) LOO
2019 Dash et al. [96] BCI 4 1635 CNN (5) train/valid/test split
2020 Dash et al. [97] BCI 8 3046
AlexNet (7)
ResNet101 (101)
Inception-ResNet-v2 (164)
train/valid/test split
2020 Dash et al. [98] BCI 8 3046 LSTM (3) N/M
2020 Yeom et al. [99] BCI 9 2160 LSTM 5Fold
2020 Dash et al. [100] BCI 4 1500 LSTM* train/valid/test split
2020 Lopopolo et van den Bosch [101] BCI 15 1377 CNN (4) houldout test set
2021 Ovchinnikova et al. [102] BCI 32 422 LF-CNN (3) nested 4Fold in a 5Fold
2022 Fan et al. [103] BCI 61 39040 CNN (3) train/test split
2018 Guo et al. [104] Event 10 102 SSAE (3) 5Fold
2019 Zheng et al. [105] Event 20 4000 CNN (8) multiple KFold
2020 Liu et al. [106] Event 20 150 CNN* KFold
2021 Zhang Chuncheng et al. [107] Event 10 N/M EEGNet (4) cross-validation
2022 Hirano et al. [108] Event 348 23177 ResNet (26) and AE 5Fold and 10Fold
2022 Bhanot et al. [109] Event 15 11000 N/M 5Fold
2022 Zhao et al. [110] Event 20 1320 DANN Leave One Out
2022 Guo et al. [111] Event 20 202 Transformer KFold
2022 Zhang et al. [112] Event 20 150
CADNet
DendriteNet
N/M
2023 Zheng et al. [68] Event 48 N/M CNN (5) 7Fold
2023 Mouches et al. [113] Event 95 40662 N/M 10Fold
2024 Wei et al. [114] Event 277 230325 N/M train/valid split
2024 He et al. [115] Event 11 35013 CNN + Attention (5) LOSO
2024 Dev et al. [116] Event 7 7060 N/M train/valid/test split
2024 Hirano et al. [117] Event 1782 60139 26 Layer scSE-ResNet 5Fold subject-wise
2016 Cichy et al. [118] Visual 15 3540 SuperVision - CNN () N/M
2017 Cichy et al. [119] Visual 15 N/M CNN (8)
Bootsrapping
resampling
2018 Seeliger et al. [120] Visual 15 1000 VGG-S - CNN (10) 10Fold
2018 Dima et al. [121] Visual 19 720
CNN (7)
AlexNet - CNN (8)
5Fold
2018 Bankson et al. [122] Visual 32 32 VGG-F - CNN (7) LOSO
2019 Rajaei et al. [123] Visual 15 1536
AlexNet - CNN (7)
HRRN (152)
LOO
2019 Kietzmann et al. [124] Visual 15 380
CNN
RCNN (6)
2Fold
bootstrapping
2020 Giari et al. [125] Visual 25 3525 AlexNet - CNN (5) N/A
2022 van Vliet et al. [126] Words 15 560 VGG (11) N/A
2023 von Seth et al. [127] Visual 36 N/M N/M N/M
2020 Donhauser et al.[128] Speech 11 77 LSTM* 7Fold
2022 Caucheteux and King [129] Auditory 92 N/M
Transformer
CNN (4, 8 or 12)
N/A
2022 Wingfield et al. [130] Speech 16 66300 MLP (6) N/A
2023 Desbordes et al. [131] Speech 11 3630
LSTM (2)
Transformer (11)
10Fold
2024 Brodbeck et al. [132]
Speech
Recognition
18 N/M N/M N/M
2024 Lyu et al. [133]
Speech
Recognition
16 5760 N/M N/M
2019 Dinh et al. [134] SL 1 N/M LSTM train/valid split
2021
Pantazis
and Adler
SL N/A N/M CNN (5) and MLP (4) N/A
2021 Dinh et al. [135] SL 1 1653 LSTM train/valid split
2022 Sun et al. [136] SL 29 620256 ResNet (4) LSTM (3) train/test split
2023 Sun et al. [137] SL 29 620256 N/M N/M
2023 O’Reilly et al. [138] SL 1 446 SimpleRNN (4) Bootstrapping
2024
Sanchez-Bort
et al
SL 1 N/M MPSS (4) N/M
2024 Jiao et al. [139] SL 1 N/M CNN + Attention N/M
2024 Yokoyama et al. [140] SL 3 N/M 4LCNN (4) train/valid sets
2016
Hyvärinen
and Morioka[141]
Preprocessing 9 N/M FC (4) M/M
2017 Garg et al. [142] Preprocessing 49 980 CNN (6) train/test split
2017 Garg et al. [143] Preprocessing 44 880 CNN (9) LOO + train/test split
2018 Croce et al. [144] Preprocessing 67 4038
CNN (6); FC (6)
CNN+FC
10Fold (trials)
2018 Hasasneh et al.[145] Preprocessing 48 1112
temporal. CNN (3)
spatial. CNN (2)
sCNN + tCNN
50Fold
2021 Feng et al. [146] Preprocessing 4 66780 GoogLeNet - CNN (144) train/valid/test split
2021 Treacher et al. [147] Preprocessing 217 294 CNN (11) 10Fold
2023 Hamdan et al. [148] Preprocessing N/M N/M AE (9) N/M
2017 Guo et al. [149] Method 3 N/M CNN (5) 10Fold
2019 Harper et al. [150] Method 72 N/M LRCN train/valid/test split
2021 Abdellaoui et al. [151] Method 18 N/M CNN + Attention train/test split
2022 Priya and Jayalakshmy [152] Method 23 13472 GoogLeNet
KFold with
holdout test set
2023 Gosti et al. [153] Method 10 N/M RHoMM - RNN (N/M) N/M
2023 Elshafei et al. [154] Method 1 166800
DNN (4)
CNN (1)
train/test split
2023 Fan et al. [155] Method 669 70000
LFCNN
VARCNN
HGRN
train/valid split
2023 Csaky et al. [156] Method 37 53100 WaveNet-based train/valid split
2023 Zhu et al. [157] Method 676 100000
MLP (1)
NICA(TCL) (3)
NICA(IIA) (3)
N/A
2024 Solana et al. [158] Method 16 N/M
ROCKET-based
models (3)
train/valid split
2024 Fan et al. [159] Method 250 29238
EEGNet
ResNet
ShuffleNet
N/M
2024 Gallard et al. [160] Method 44 N/M CycleGAN N/A
2024 Chou et al. [161] Method 17 N/M CNN (3) train/valid/test split
2024 Csaky et al. [162] Method 15 3540 GPT train/valid/test split
2024 Gideoni et al. [163] Method 6 N/M MLP and CNN train/valid/test split
2024 Ferrante et al. [164] Method 4 N/M CNN and CLIP N/A
2.2 Types of studies
In conducting this review, we opted to categorize the included studies into three distinct groups: (1) ’Classification’, (2) ’Modeling’, and (3) ’Other’. The ’Classification’ category was split into four sub-groups: decoding, brain-computer interfaces (BCI), clinical applications or computer-aided diagnosis (CAD), and event detection. Studies under ’decoding’ primarily leverage classification as a statistical tool to identify the involvement of specific brain features in certain tasks. The BCI subcategory includes studies that incorporate ANNs in their BCI framework. Meanwhile, the ’clinical’ subcategory includes studies that utilize ANNs to enhance clinical diagnosis or prognosis, and the ’event detection’ subcategory regroups studies using ANNs to continuously detect events in MEG data. Studies in the ’Modeling’ category primarily focus on comparing activations across ANNs and the brain. The ’Other’ category gathers the studies that do not fit the aforementioned groups and includes research utilizing ANNs for preprocessing tasks such as artifact detection and removal or source localization. The distribution across more specific subcategories is given in 6 and illustrated in figure 2.
Category Subcategory Study
Classification Decoding [56], [64], [63], [57], [61], [53], [60], [52], [51], [55], [77]
[62], [50], [49], [58], [70], [54], [74], [65], [71],
[72], [59], [68], [73], [69], [66], [67], [75], [76]
BCI [91], [97], [100], [98], [96], [92], [94], [93],
[103], [102], [99], [95], [101], [115]
Clinical [79], [78], [83], [85], [86], [81], [80], [84], [82], [89],
[90], [87], [88], [112], [110], [109], [108]
Event detection [165], [105], [111], [106], [107], [104], [113], [114], [116][117]
Modeling Visual cortex [121], [122], [118], [119], [124], [120], [125], [126] [123], [127]
Auditory cortex [128], [129], [131], [130], [132], [133]
Other Preprocessing [144], [143], [146], [166], [142], [141], [147], [148],
Methods [156], [151], [149], [150], [152], [155], [153], [160],
[154], [162], [158], [159], [161], [164], [163], [157]
Source localization [134], [167], [136], [138], [135], [137], [168],
[139], [164]
@@ 2.3 Notes on terminology
Some of the ML nomenclature overlaps with terminology used in M/EEG research, which can sometimes lead to confusion. In particular, the term ’epoch’ in deep learning generally describes a complete pass of the training data, where the entire dataset is used once to update the model parameters. Meanwhile, in the field of M/EEG, it usually describes a segment of data (or a single trial). In addition, the term ’samples’ in ML usually refers to individual examples from the dataset, which should not be confused with the use of the same term to refer to a data point in a M/EEG time series, where the parameter known as ’sampling rate’ (Hz) refers to the rate at which the data is assessed, i.e., data points per second.
@@ 3 Results
The methodology explained in the previous section (2) led to the identification of 119 relevant studies. Among the papers considered, 102 were peer-reviewed, while 17 had not yet undergone peer review at the time of writing. Our categorization method resulted in the distribution of 70 papers under ’Classification’, accounting for 58.8% of the total; 16 papers were categorized as ’Modeling’, making up 13.4%; and 33 papers fell into the ’Other’ category, representing 27.8% of the study corpus included in this review (See figure 2). The category containing most of the papers is the ’Classification’ category, followed by ’Other’; the category with the least amount of studies is ’Modeling’. Generally, the trend of using ANNs with MEG data has increased over the years across all categories, as depicted in figure 2. A similar growth trajectory is anticipated for studies on this topic, analogous to the rapid expansion of research utilizing ANNs for EEG data analysis ([5]).
@@
3.1 General overview
The in-depth survey revealed several key observations about this emerging field (figure 2). Interestingly, out of the 119 selected studies for this review, 65 use less than 20 subjects, 35 use between 20 and 100 subjects, and 16 use more than 100 subjects. The remaining studies use synthetic MEG data or do not mention the number of subjects, or patients included in the study. The number of samples (number of trials or epochs) in datasets varies across a wide range, from 17 to 2304000 (see figure 2), with an average of 102247, a standard deviation of 383501, and a median of 4000. However, a few outliers significantly impact the average and standard deviation of the number of samples used in the included studies (Correcting for outliers gives an average sample number of 13246 and a standard deviation of 21147). Furthermore, the length of the data segments fed to the neural network varies between 300ms to 10s. The most commonly used sampling frequencies are 1000 Hz (26 papers) and 250 Hz (17 papers) (as illustrated by figure 3), and the median across all studies was 250 Hz.
Figure 3: Distribution of sampling frequencies (in Hz) reported across the reviewed studies (N=110 reporting). The height of each bar indicates the total number of studies using a specific frequency. Stacked colors represent the main study categories (yellow: ’Other’, red: ’Modeling’, blue: ’Classification’). The most frequently reported rates were 1000 Hz (N=26) and 250 Hz (N=17).